OpenStudy (anonymous):

Find the two products below. Compare and contrast, in complete sentences, the similarities and differences of the two. (x + 6)(x − 6) and (x − 6)(x − 6)

OpenStudy (anonymous):

(a-b)(a+b) = a^2 - b^2

OpenStudy (anonymous):

\[(a -b)(a+b) = a^2 - b^2\] \[(a-b)(a-b) = (a-b)^2\] Use them..

OpenStudy (anonymous):

\[(a-b)(a-b) = (a-b)^2 = a^2 + b^2 - 2ab\]

OpenStudy (anonymous):

[x+6][x-6] is a difference between squares [x-6][x-6] will produce a perfect trinomial square when multiplied. this means it has double roots. [x-6][x-6]=[x-6]^2= double roots

OpenStudy (anonymous):